Five years ago, Sean was three times as old as his brother. Today Sean is twice as old as his brother. How many years older than his brother is Sean? How old is Sean now?
Answer by Earlsdon(6294)

Let S = Sean's present age and B = his brother's present age, then:
1) S-5 = 3(B-5) "Five years ago, Sean was three times as old as his brother (was five years ago)."
2) S = 2B "Today Sean is twice as old as his brother."
Substitute equation 2) into equation 1) to get:
(2B)-5 = 3(B-5) Simplify.
2B-5 = 3B-15 Subtract 2B from both sides.
-5 = B-15 Now add 15 to both sides.
10 = B and S = 2(10) = 20.
Sean is 20 years old and his brother is 10 years old, so Sean is 10 years older than his brother.